Nerissa has the option of either paying her obligation of P10,500 today or settling it for P11,200 after 6 months. Find the simple interest rate at which she would be charged?
Answer by math_tutor2020(3817)


i = 11,200 - 10,500 = 700 is the amount of interest

P = 10500 is the amount loaned

t = 6 months = 6/12 = 0.5 year is the time value

Simple interest formula
i = P*r*t
700 = 10500*r*0.5
700 = 10500*0.5*r
700 = 5250*r
r = 700/5250
r = 0.1333

Answer: Roughly 13.33%